Solve: z(x+2y)p-z(2x+y)q=y^2-x^2


1
Expert's answer
2021-11-17T14:46:58-0500

Solution;

"z(x+2y)p-z(2x+y)q=y^2-x^2"

The auxiliary equation are;

"\frac{dx}{z(x+2y)}=\frac{dy}{-z(2x+y)}=\frac{dz}{y^2-x^2}"

Considering the first two equations;

"\frac{dx-dy}{3z(x+y)}=0" .....(1)

"dx=dy"

Integrating;

"x=y+c_1"

Considering (1) and the last auxillary equation;

"\frac{dx-dy}{3z(x+y)}=\frac{dz}{x^2-y^2}"

"(x-y)(dx-dy)=3zdz"

"xdx-xdy-ydx+ydy=3zdz"

Integrating;

"\frac{x^2}{2}-2xy+\frac{y^2}{2}=\frac 32z^2+c_2"

"x^2-4xy+y^2=3z^2+c_2"

The solution of the equation is;

"F(x-y,x^2+y^2-3z^2-4xy)=0"